Seven years ago a father was 11 times the age of his daughter. Now, he is 4 times his daughters age. What is his age? I have the answer, just can't put a formula to it. Need the formula please.
Thank you!

Answer by ewatrrr(24785)


Hi
Now, he is 4 times his daughters age
Let x and 4x represent daughter and father's ages NOW
question States***
4x - 7 = 11(x-7) |seven years ago...
solving for x
4x - 7 = 11x - 77
70 = 7x
x = 10, daughter's age now. Father is now 40. (4*10)
CHECKING our Answer
33 = 11*3